OpenStudy (anonymous):

how do I find the n th term in the following sequence? 1,3,6,10,15,21... a30

OpenStudy (anonymous):

we seem to be doing the same hw :D

OpenStudy (anonymous):

an=a1+(n−1)d

OpenStudy (anonymous):

isn't that when it is constant?

OpenStudy (anonymous):

yeah, i didn't see that....thinking ;/

OpenStudy (anonymous):

did you try the geometric sequence formula?

OpenStudy (anonymous):

\[a _{n}=a _{1}r ^{n-1}\] ?

OpenStudy (anonymous):

this is 1 1+2 1+2+3 1+2+3+4 1+2+3+4+5

OpenStudy (anonymous):

formula is \[\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and so \[a_{30}=\frac{(30)(31)}{2}=15\times 31=465\]

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (anonymous):

yw